OpenStudy (study312):

The table below shows the proportional relationship between the weight of cereal, in ounces, and the number of packs of cereal: Number of Packs of Cereal Number of Ounces of Cereal 2 34 3 51 4 68 What is the constant of proportionality?

Mehek (mehek14):

\(\dfrac{51-34}{3-2}\)

OpenStudy (cal.lavender):

Is there more @Study312 ?

OpenStudy (study312):

Number of packs/ number of ounces 2 / 34 3 / 51 4 / 68

OpenStudy (study312):

that's to look at it better

OpenStudy (study312):

so why do I subtract 34 from 51?

Mehek (mehek14):

constant of proportionality = slope

Mehek (mehek14):

so when you use the slope formula, \(\dfrac{51-34}{3-2}\)

OpenStudy (study312):

so 51 - 34 = 17 3 - 2 = 1 17/1

Mehek (mehek14):

so the constant of proportionality is 17

OpenStudy (study312):

ok thanks

Mehek (mehek14):

yw :)